OpenStudy (anonymous):

What is the fractional notation of 0.92 when 92 is underlined

OpenStudy (anonymous):

I'm not sure what "when 92 is underlined" means. But the fractional notation of 0.92 is 92/100 = 23/25

OpenStudy (anonymous):

What would 3.092 be in fractional notation?

OpenStudy (anonymous):

Do you mean 0.929292...92?

OpenStudy (anonymous):

yes larry!

OpenStudy (anonymous):

0.92 repeating = (92/99)

OpenStudy (reemii):

a = 0.929292... 100a = 0.9292... * 100 = 92.929292.... remove 0.929292... from it. 100a - a = 92 100a - a = 99a also. larry gave the answer

OpenStudy (anonymous):

Okay, what about 3.092 then

OpenStudy (reemii):

3.092929292... ?

OpenStudy (anonymous):

As an integer or simplified fraction

OpenStudy (reemii):

ok. same technique. the repeating pattern is "92", the length of it = 2. so you'll start by multiply it by 100. a = 3.0929292... 100a = 309.29292... 99a = 309.292929... - 3.09292929... = 309.29 - 3.09 = 306.2 almost done!

OpenStudy (reemii):

can you go on?

OpenStudy (anonymous):

No, I'm not sure. I thought 306.2 was the answer